# Modelling two-dimensional droplet rebound off deep fluid baths

### v2

@@ 4 Results
We present the numerical results of the lubrication-mediated quasi-potential model, and begin by introducing a single test case which sits well within a stable parameter regime, using this to discuss the nuances of the model and how it compares to the DNS results. We proceed through a parameter sweep, highlighting the range of the model and general trends of droplet trajectories observed within a two-dimensional rebound case, all of which we present in the absence of gravity, as it is difficult for the droplet to detach under the effects of gravity in the two dimensional system. Finally, we discuss the extensions to more complicated rebounds that can also be obtained, such as sustained periodic bouncing behaviour in the presence of Faraday forcing.
@@
Figure 2: Comparison of lubrication-mediated model for the solid and liquid drop cases and DNS results for a droplet with $R_{0}=0.2\times 10^{-3}$ m. released with velocity $W_{0}=-0.2$ms-1 from a centre of mass height of $2R_{0}$. Shown are the centre of mass and poles of both droplets, and the free surface directly beneath the south pole. Simulations are run until the centre of mass returns to a height $R_{0}$.
4.1 Model output
@@
We first consider a droplet with radius $R_{0}=0.2\times 10^{-3}$ m and initial velocity of $W_{0}=-0.2$m/s, and compare the cases where the droplet is either solid or liquid to highlight the effect of droplet deformation on the rebound behaviour. In both cases we take the liquid in the bath to be water, and we take the droplet to be water or solid with the density of water. Parameter values are presented in Appendix A. To consider solid droplets, the surface of the sphere is fixed and not permitted to deform setting $c_{n}=0$, hence $\eta_{d}(\theta,t)=R_{0}$.
@@
Figure 2 presents the trajectories in the solid and deformable lubrication-mediated models, the latter also being contrasted to their counterpart DNS results. The lubrication-mediated rebound in the deformable case shows good agreement with DNS on the whole, though has a much smoother contact trajectory, which can be explained by the enhanced viscosity used during impact causing it to behave more closely to the solid drop during contact. Post contact, the droplet poles in both computations show oscillation generated by the deformation during contact, however the DNS oscillations are larger. The solid case has a slower exit velocity, due to the solid drop losing more kinetic energy to the bath, and not storing energy during impact as we shall see below. Henceforth, results are plotted nondimensionally using a capillary timescale $t_{\sigma}=\sqrt{\rho R_{0}^{3}/\sigma}$ and pressure $\sigma/R_{0}$ and with radius and initial velocity being used for typical length scale and velocity. We discuss the choice of enhanced viscous damping briefly in Appendix B.
@@
Figure 3: Energy balance for the solid and deformable case of droplet impacts. The energy in the system starts in the droplet kinetic energy $K\!E_{\text{drop}}$, and is transferred to the deformation and waves on the bath, and, in the deformable case, to the droplet oscillations during impact. The system energy is computed to be the sum of these terms, from which we can the loss of system energy due to the dissipation within the system.
Once impact begins energy is transferred to the free surface motion and (for deformable drops) to droplet deformation. The effect of deformation on the overall dynamics is clear: deformation lessens the energy transferred into the waves of the bath, instead transferring energy to the oscillations within the droplet. Some of the droplet energy is returned in a rebounding effect and contributes to the upwards velocity of the droplet, increasing its kinetic energy relative to the solid case. This results in the higher exit velocity seen in the deformable case despite a lower final total energy. The effect of viscous damping is seen through the loss of total energy which is enhanced in the deformable drop case. The Ohnesorge number appears naturally as a decay constant, with a typical decay over a capillary timescale of $e^{-(2\pi)Oh}\approx 0.95$, matching qualitatively the computational results.
@@
Figure 4: Droplet deformation and droplet rebound.
@@
Figure 5: Air layer height for the lubrication-mediated model (left) and DNS (right). The height is plotted dimensionally. At the edges of the lubrication region and during the pinch before liftoff the film thicknesses may be as small as $5\times 10^{-7}$ m.
A key distinction between the work presented here and the current literature is the dynamic inclusion of the lubrication layer within the model. In Figure 5 we show a spatio-temporal plot of the lubrication layer height, over $\Omega_{l^{*}}(t)$. There are three main stages of rebound. The first stage happens quickly as the droplet “impacts” the bath and the lubrication region spreads out on timescales shorter than $t_{\sigma}$. Rapidly, the lubrication region reaches its maximum extent and in the second stage begins to recede as the droplet peels away from the bath. During this stage, a narrow region forms near the outer edge where the layer depth is much thinner than at the centre of the lubrication region. If coalescence were permitted by the model, this would be the point (or ring) of first contact that is described in the literature (Thoroddsen et al., 2003). We could expect to see contact when this depth is sufficiently small for atomic scale effects (e.g. Van der Waals forces) to become relevant. The third stage is pinchoff, which happens on a very short timescale and with negative pressure (suction) near the pichoff point.
Figure 6: Dimensionless pressure $P^{*}$ within the air layer as determined by equation (48), presented as (a) a spatio-temporal plot, and (b) at different times. The large spike of pressure at initial “impact” quickly spreads out and retains a flattened profile for most of contact. As the pressure diminishes in the later stages a region of low pressure forms on the boundary and propagates inwards, generating a suction effect. Eventually the suction region reaches the centre of the droplet, before detachment. The black lines in panel (a) correspond to the snapshots within figure 7.
@@
Figure 7: Four representative snapshots of nondimensional layer height $h^{*}$, pressure $P^{*}$ and flux $Q^{*}$ within the lubrication region $\Omega_{l^{*}}(t)$. The times correspond to solid lines on Figure 6(a), chosen to highlight the change in behaviour between at distinct time regions during impact; initial contact (stage 1), widest point of contact, slow detachment (stage 2), and pinch-off (stage 3). Note that the axes change between each snapshot.
The centre lubrication region in both the lubrication-mediated model and DNS presented is of near constant thickness of $\mathcal{O}(1)\mu$ and pressure $\mathcal{O}(\sigma/R_{0})$ for most of the impact, a quasi-static air cushion. This was experimentally observed in Tang et al. (2019), and corroborated in DNS computations Galeano-Rios et al. (2021); Alventosa et al. (2023); Sprittles (2024). The region where the layer pinches at its boundaries lines up with the maxima (at the transition between stage 1 and stage 2) and minima (during stage 2) at at the edge of the pressure profile.
@@
Figure 7(b) corresponds to the widest the contact region, which occurs before the drop has reached maximum depth. The pressure is near constant (and hence flux is near zero) for the majority of the lubrication region with a very slight maximum near its edge and a smooth decay to atmospheric pressure at the outer region of the layer. The maximum is more pronounced in three-dimensions (Phillips & Milewski, 2024). The narrowing of the layer at its edge is now apparent.
Figure 7(c) corresponds to stage 2. The layer edge is propagating inwards towards the centre of the drop, the trapped air region contracts and the narrowing is more pronounced. There is a suction zone (negative pressure) and corresponding (negative) flux of outside air towards the the outer boundary of the layer.
@@ 4.2 Parameter Study
Figure 8: The evolution of four runs from point of impact to detachment. The top two runs correspond to a smaller droplet of radius $R_{0}=0.5\times 10^{-4}$m, and the lower two correspond to a larger droplet with radius $R_{0}=0.7\times{10^{-3}}$.
@@
Figure 9: Comparisons of Coefficient of restitution $\alpha$, penetration depth $\delta$ and contact time $t_{c}$, to Weber number (a,c,e) and Ohnesorge number (b,d,e) for a parameter sweep of values. The dimensionless parameter not plotted is indicated by the colour gradient on the opposite graph. Lubrication-mediated numerical results are presented by circles ($\bullet$), and triangles are used to indicate DNS results ($\triangle$) which correspond to a subset of the deformable parameter cases. That subset is indicated by the circles being filled.
Figure 10: $\alpha$, $\delta$, $\tau$ plotted against Reynolds number $Re$. The colours correspond to varying Weber numbers, with connected lines corresponding to fixed Ohnesorge numbers. Here we have included the solid model results by the squares ($\square$), and as before circles ($\bullet$) denote the deformable model results, with their DNS counterpart results shown with triangles ($\triangle$).
Figures 9 and 10 displays the three parameters of interest $\alpha$, $\delta$, $\tau$, against $W\!e$, $Oh$, $Re$. We include solid and drop simulations for the lubrication-mediated model, and drop simulations for DNS.
@@
For the deformable droplet, as shown in Figure 9, the coefficient of restitution generally decreases with $Oh$ and $W\!e$ (with the other parameter fixed). An exception to this is the larger Ohnesorge cases (corresponding to smaller droplets) where there is a local maximum in $\alpha$ as a function of $W\!e$. This has been observed in DNS and kinematic match modelling of three dimensional droplets (Alventosa et al., 2023). Their work is for axisymmetric 3D simulations (without resolving a lubrication layer in the kinematic match model) and therefore not directly comparable to ours. Nevertheless generally trends are preserved, despite the restitution coefficients in 3D being about a factor of 2 larger. Our lubrication mediated results and trends generally corroborate with the DNS results herein, though DNS results are monotonic and have slightly higher coefficients of restitution, particularly at lowest $Oh$ and intermediate $W\!e$.
Regarding maximum depth, as shown in Figure 9, there is agreement on trends between the model and DNS. Although the lubrication-mediated model always reached a lower depth than the DNS model, particularly at higher $W\!e$, both models show that penetration depth is near independent from $Oh$ and has a strong positive correlation with $W\!e$. We note that a similar discrepancy between DNS and the kinematic match model was observed in Alventosa et al. (2023) and that the discrepancy fell within the experimental range. As shown in Figure 10, the difference in maximum depth between the solid and deformable lubrication-mediated results was small.
@@
5 Extensions and Discussion
@@
In the previous section we presented a series of results obtainable by the lubrication mediated model and highlighted their general agreement with well validated DNS methods through a parameter study. We will now highlight the robustness of the lubrication mediated model through demonstrating sustained bouncing behaviour under the influence of gravity and a vertical oscillation of the bath. This justifies the potential of the model to be used within further studies of Faraday pilot-wave models: in particular tunnelling studies (Nachbin et al., 2017) use two dimensional droplets. Finally we will conclude this paper with a discussion of the lubrication-mediated model.
5.1 Repeated Bouncing
@@
Thus far we have reviewed the results of the lubrication-mediated model of the two-dimensional droplet rebounding off a deep liquid bath. A natural extension to this system is the inclusion of repeated bounces with the inclusion of other effects. With the addition of gravity, after the initial rebound and detachment, the droplet undergoes ballistic motion and falls back onto the bath. As noted by Couder et al. (2005), a way to investigate repeated bouncing behaviour of droplets is to vibrate the liquid bath and achieve a sustained periodic behaviour for a vibrational forcing of sufficient amplitude. Both gravity and vibrational forcing are straightforward additions to the model and a sinusoidal forcing in the gravity term is introduced with the functional form $G(t)=g(1-\Gamma\cos(\omega_{f}t))$. For sufficiently large $\Gamma$ such a forcing generates the Faraday instability: growing standing waves on the free surface known as Faraday waves. The threshold at which Faraday waves are excited is denoted $\Gamma_{F}$. In order for sustained bouncing behaviour to occur, the bath is instead oscillated close to, but below this threshold, that is, for some $\Gamma<\Gamma_{F}$. In what follows we present the results when such a forcing is introduced to the lubrication-mediated model.
@@
Following the methodology presented in Milewski et al. (2015), we incorporate the forced oscillations of the bath as an oscillating frame of reference. For a given oscillation frequency $\omega_{f}$ and forcing amplitude $\Gamma$, the equations derived in Section 2 for the liquid bath are now given as
@@
and the equations governing the droplet motion are
@@
Figure 11 demonstrates the periodic bouncing behaviour obtained with the lubrication-mediated model. To capture these results, the numerical domain was shortened to an integer ratio of the Faraday wavelength of the bath, specifically, corresponding to $L=12$cm. The droplet was then released from a small height, with an initial velocity of $0.1$ms-1. The choice of initial velocity is unimportant as long as the droplet is able to rebound, as the system will eventually stabilise to a periodic state with these parameters.
Figure 11: Periodic droplet rebounds obtained for viscous silicone oil with radius $R=8.3\times 10^{-4}$m, on a bath oscillated at $80$Hz with forcing amplitude $\Gamma=3$. Plotted with the oscillation period of an undisturbed free surface behind to highlight deviation away from this rest state. The dynamics settles on a (4,2) mode with period $0.05$s.
As shown in Figure 11 the stable state is a $(4,2)$ mode (Milewski et al., 2015), which is denoted as such due to the periodic motion requiring 4 oscillations of the bath for every 2 rebounds of the droplet. In the figure, the dashed line indicates the unperturbed surface, as though the behaviour of the droplet were omitted from the system. The upper line then represents the south pole of the drop, and the lower blue line corresponds to the free surface of the liquid bath.
@@
5.2 Conclusion

### v3

@@ 4 Results
We present the numerical results of the lubrication-mediated quasi-potential model for both solid and deformable droplets, and begin by introducing a single test case which sits well within a stable parameter regime, using this to discuss the nuances of the model and how it compares to the DNS results in the deformable scenario. We proceed through a parameter sweep, highlighting the range of the model and general trends of droplet trajectories observed within a two-dimensional rebound case, all of which we present in the absence of gravity, as it is difficult for the droplet to detach under the effects of gravity in the two-dimensional system due to the stronger relative influence in the infinite-cylinder regime than a true spherical case [42]. Finally, we discuss the extensions to more complicated rebounds that can also be obtained, such as sustained periodic bouncing behaviour in the presence of Faraday forcing.
@@
Figure 2: Comparison of lubrication-mediated model for the solid and liquid drop cases, and DNS results for a deformable droplet, with initial undeformed radius $R_{0}=0.2$mm released with velocity $W_{0}=-0.2$ms-1 from a centre of mass height of $2R_{0}$. Shown are the centre of mass and poles of both droplets, and the free surface directly beneath the south pole. Simulations are run until the centre of mass returns to a height $R_{0}$.The liquid drop has the physical properties of water, and the solid impactor has the same density, with all other parameters indicated in Table 1.
4.1 LM model investigation
@@
Within Figures 2-7 we first consider a droplet with radius $R_{0}=0.2$ mm and initial velocity of $W_{0}=-0.2$ms-1, and compare the cases where the droplet is either solid or liquid to highlight the effect of droplet deformation on the rebound behaviour. In both cases we take the liquid in the bath to be water, and we take the droplet to be water or solid with the density of water. Parameter values are presented in Appendix A. To consider solid droplets, the surface of the sphere is fixed and not permitted to deform setting $c_{n}=0$, hence $\eta_{d}(\theta,t)=R_{0}$. For further discussion of how viscous damping during impact affects rebound dynamics (including consideration of the soft solid limit), please see Appendix C.
@@
Figure 2 presents the trajectories in the solid and deformable lubrication-mediated models, the latter also being contrasted to their counterpart DNS results. The lubrication-mediated rebound in the deformable case shows good agreement with DNS on the whole, though has a much smoother contact trajectory, which can be explained by the enhanced viscosity used during impact causing it to behave more closely to the solid drop during contact. Post contact, the droplet poles in both computations show oscillation generated by the deformation during contact, however the DNS oscillations are larger. The solid case has a slower exit velocity, due to the solid drop losing more kinetic energy to the bath, and not storing energy during impact as we shall see below. Henceforth, results are plotted nondimensionally using a capillary timescale $t_{\sigma}=\sqrt{\rho R_{0}^{3}/\sigma}$ and pressure $\sigma/R_{0}$ and with radius and initial velocity being used for typical length scale and velocity. We discuss the choice of enhanced viscous damping briefly in Appendix C.
@@
Figure 3: Energy balance for the solid and deformable impactors for the lubrication-mediated model, with parameters as shown in Figure 2, and in particular initial radius $R_{0}=0.2$mm, and initial velocity $W_{0}=-0.2$ms-1. The energy in the system starts in the droplet kinetic energy $K\!E_{\text{drop}}$, and is transferred to the deformation and waves on the bath, and, in the deformable case, to the droplet oscillations during impact. The system energy is computed to be the sum of these terms, from which we can the loss of system energy due to the dissipation within the system.
Once impact begins energy is transferred to the free surface motion and (for deformable drops) to droplet deformation. The effect of drop deformation on the overall dynamics is clear: deformation lessens the energy transferred into the waves of the bath, instead transferring energy to the oscillations within the droplet. Some of the droplet energy is returned in a rebounding effect and contributes to the upwards velocity of the droplet, increasing its kinetic energy relative to the solid case. This results in the higher exit velocity seen in the deformable case despite a lower final total energy. The effect of viscous damping is seen through the loss of total energy which is enhanced in the deformable drop case. The Ohnesorge number appears naturally as a decay constant, with a typical decay over a capillary timescale of $\exp{(-2\pi Oh)}\approx 0.95$, matching qualitatively the computational results.
@@
Figure 4: Droplet deformation and droplet rebound characteristics for the deformable droplet lubrication-mediated model, as shown in Figure 2, with initial radius $R_{0}=0.2$mm, and initial velocity $W_{0}=-0.2$ms-1.
@@
Figure 5: Dimensionless air layer height for a deformable droplet impact for the lubrication-mediated model (left) and DNS (right) as shown in Figure 2, with initial radius $R_{0}=0.2$mm, and initial velocity $W_{0}=-0.2$ms-1. The time axis in both figures is restricted to the approximate contact time, which begins from time $t/t_{\sigma}\approx 3$, to highlight impact behaviour. At the edges of the lubrication region and during the pinch before lift-off the dimensional film thicknesses may be as small as $0.5\;\mu$m.
A key distinction between the work presented here and the current literature is the dynamic inclusion of the lubrication layer within the model. In Figure 5 we show a spatio-temporal plot of the lubrication layer height, over $\Omega_{l^{*}}(t)$. There are three main stages of rebound. The first stage happens quickly as the droplet “impacts” the bath and the lubrication region spreads out on timescales shorter than $t_{\sigma}$. Rapidly, the lubrication region reaches its maximum extent and in the second stage begins to recede as the droplet peels away from the bath. During this stage, a narrow region forms near the outer edge where the layer depth is much thinner than at the centre of the lubrication region. If coalescence were permitted by the model, this would be the point of first contact that is described in the literature [43]. We could expect to see contact when this depth is sufficiently small for atomic scale effects (e.g. Van der Waals forces) to become relevant. The third stage is pinchoff, which happens on a very short timescale and with negative pressure (suction) near the pinchoff point.
Figure 6: Dimensionless pressure $P^{*}$ within the air layer as determined by equation (48) for the deformable lubrication-mediated rebound presented in Figure 2, with initial radius $R_{0}=0.2$ mm and initial velocity $W_{0}=-0.2$ms-1, presented as (a) a spatio-temporal plot (with impact beginning from time $t/t_{\sigma}\approx 3$), and (b) at different times. The large spike of pressure at initial “impact” quickly spreads out and retains a flattened profile for most of contact. As the pressure diminishes in the later stages a region of low pressure forms on the boundary and propagates inwards, generating a suction effect. Eventually the suction region reaches the centre of the droplet, before detachment. The black lines in panel (a) correspond to the snapshots within figure 7.
@@
Figure 7: Four representative snapshots of nondimensional layer height $h^{*}$, pressure $P^{*}$ and flux $Q^{*}$ within the lubrication region $\Omega_{l^{*}}(t)$ during the deformable lubrication-mediated rebound presented in Figure 2 with initial radius $R_{0}=0.2$mm and velocity $W_{0}=-0.2$ms-1. The times correspond to solid lines on Figure 6(a), chosen to highlight the change in behaviour between at distinct time regions during impact; initial contact (stage 1), widest point of contact, slow detachment (stage 2), and pinch-off (stage 3). Note that the axes change between each snapshot.
The centre lubrication region in both the lubrication-mediated model and DNS presented is of near constant thickness of $\mathcal{O}(1)\ \mu$m and pressure $\mathcal{O}(\sigma/R_{0})$ for most of the impact, a quasi-static air cushion. This was experimentally observed in Tang et al [40], and corroborated in DNS computations [29, 33, 5]. The region where the layer pinches at its boundaries lines up with the maxima (at the transition between stage 1 and stage 2) and minima (during stage 2) at the edge of the pressure profile.
@@
Figure 7(b) corresponds to the widest contact region, which occurs before the drop has reached maximum depth. The pressure is near constant (and hence flux is near zero) for the majority of the lubrication region with a very slight maximum near its edge and a smooth decay to atmospheric pressure at the outer region of the layer. The maximum is more pronounced in three-dimensions [44]. The narrowing of the layer at its edge is now apparent.
Figure 7(c) corresponds to stage 2. The layer edge is propagating inwards towards the centre of the drop, the trapped air region contracts and the narrowing is more pronounced. There is a suction zone (negative pressure) and corresponding (negative) flux of outside air towards the outer boundary of the layer.
@@ 4.2 Parameter study
Figure 8: The evolution of four parameter regimes for the deformable LM model from point of impact to detachment. The top two runs correspond to a smaller droplet of radius $R_{0}=0.05$mm, and the lower two correspond to a larger droplet with radius $R_{0}=0.7$mm. The Weber number was varied by changing the initial velocity of the droplets respectively.
@@
Figure 9: Comparisons of Coefficient of restitution $\alpha$, penetration depth $\delta$ and dimensionless contact time $\tau$, to Weber number (a,c,e) and Ohnesorge number (b,d,e) for a parameter sweep of values. The dimensionless parameter not plotted is indicated by the colour gradient on the opposite graph. Lubrication-mediated numerical results are presented by circles ($\bullet$), and triangles are used to indicate DNS results ($\triangle$) which correspond to a subset of the deformable parameter cases. The respective model subset is indicated by the circles being filled.
Figure 10: Coefficient of restitution $\alpha$, penetration depth $\delta$ and dimensionless contact time $\tau$ plotted against Reynolds number $Re$. The colours correspond to varying Weber numbers $W\!e$, with connected lines corresponding to fixed Ohnesorge numbers $Oh$. Here we have included the solid model results by the squares ($\square$), and as before circles ($\bullet$) denote the deformable model results, with solid lines highlighting their DNS counterpart results shown with triangles ($\blacktriangle$).
Figures 9 and 10 display the three parameters of interest $\alpha$, $\delta$, $\tau$, against $W\!e$, $Oh$, $Re$. We include solid and drop simulations for the lubrication-mediated model, and drop simulations for DNS.
@@
For the deformable droplet, as shown in Figure 9, the coefficient of restitution generally decreases with $Oh$ and $W\!e$ (with the other parameter fixed). Exception to this are the larger Ohnesorge cases (corresponding to smaller droplets) where there is a local maximum in $\alpha$ as a function of $W\!e$. This has been observed in DNS and kinematic match modelling of three-dimensional droplets [33]. In the respective work axisymmetric 3D simulations (without resolving a lubrication layer by using a kinematic match model) are compared to other studies, including experimental observations. Broadly, between the three-dimensional data therein and our two-dimensional computations, general trends are preserved despite the restitution coefficients in three dimensions being approximately a factor of two larger than in two dimensions. Our lubrication-mediated two-dimensional model results and trends corroborate far better with the two-dimensional DNS results herein, the only differences being that DNS results are monotonic in $W\!e$ and have slightly higher coefficients of restitution, particularly at lowest $Oh$ and intermediate $W\!e$.
Maximum depth results, as shown in Figure 9, indicate strong agreement on trends between the model and DNS. Although the lubrication-mediated model always reached a lower depth than the DNS results, particularly at higher $W\!e$, both models show that penetration depth is nearly independent of $Oh$ and has a strong positive correlation with $W\!e$. We note that a similar discrepancy between DNS and the kinematic match model was observed in Alventosa et al [33] and that the discrepancy fell within the experimental range. As shown in Figure 10, the difference in maximum depth between the solid and deformable lubrication-mediated results was small.
@@
Overall the agreement between the LM model and the DNS is highly encouraging, with key trends replicated successfully and explainable discrepancies owing to modelling assumptions or anticipated numerical effects on both sides. More specifically, DNS results can be summarised as exhibiting stronger pool deformation on impact, larger coefficients of restitution and a wider variation interval in contact times than their reduced-order model counterparts. Figure 14 in Appendix C encapsulates some of the above features for a typical comparison case. The viscosity enhancement in the model provides a useful handle for navigating the deformable to non-deformable impactor property set, while addressing numerical challenges in terms of stability in its discretisation scheme. The lower the value of this parameter (down to a tractable threshold level), the closer the results become to the DNS benchmark liquid impact calculation, with pool deformation depth converging towards DNS-obtained results for the lowest enhanced viscosity values used in the model in the liquid impact case. The scenarios described by more pronounced pool deformation naturally lead to stronger bouncing, expressed in terms of larger coefficients of restitution, with any differences arguably becoming more prominent as $W\!e$ increases and fully nonlinear DNS calculations are expected to start diverging from the simplified model counterpart results. The behaviour of the impactor during the spreading and rebound stages in the two approaches is another point of differentiation, with more nonlinear behaviour in moderate to large deformation cases in the DNS datasets visible even as part of earlier air layer height dynamics studies, as shown for example in Figure 5 in terms of the radial extent of the contact region. More deformed droplets on impact in the DNS lead to larger contact radii and marginally thinner air layers, while the rebound phase is also slightly accelerated by comparison, correlating with the earlier observations on the small coefficient of restitution differences between the two employed methodologies.
We postulate that the impact depth $\delta$ depends little on $Oh$ because the maximum depth is achieved early in the impact cycle. Thus most energy loss up to this stage is due to the transfer of energy to motion in the bath and not to viscous effects. One might expect $\delta\sim\sqrt{W\!e}$ from a model of a linear spring whose initial condition is proportional to the impact velocity $W_{0}$. The contact time and coefficient of restitution, however, depend more strongly on viscosity. When the data is plotted against $\sqrt{W\!e}\,Oh$ in Figure 11 we observe a better description of the behaviour and collapse of the data. A justification for this scaling arises from postulating an exponential decay rate for a linear spring model with decay rate proportional to $Oh$. This results in a linearly approximated viscous loss proportional to $Oh$, for small $Oh$ and finite time. For the contact times, we note that some high Weber number DNS data fall above lower Weber number data in this scaling. We conjecture that this is due to the increase in the (integer) number of oscillations of the droplet during impact in the DNS. These additional oscillations are not present in the LM model due to the additional damping effects introduced there.
Figure 11: Collapsed data for coefficient of restitution $\alpha$ and dimensionless contact time $\tau$, both plotted against $\sqrt{W\!e}\,Oh$. The colour gradient corresponds to Weber number $W\!e$. The circles ()($\bullet$)( ∙ ) denote deformable LM model results, and DNS data is shown using triangles ()($\blacktriangle$)( ▲ ).
5 Repeated bouncing
@@
In the previous section we presented a series of results obtainable by the lubrication mediated model and highlighted their general agreement with well validated DNS methods through a parameter study. We will now highlight the robustness of the lubrication mediated model through demonstrating sustained bouncing behaviour under the influence of gravity and a vertical oscillation of the bath. This justifies the potential of the model to be used within further studies of Faraday pilot-wave models: in particular tunnelling studies [19] use two-dimensional droplets.
@@
Thus far we have reviewed the results of the lubrication-mediated model of the two-dimensional droplet rebounding off a deep liquid bath. A natural extension to this system is the inclusion of repeated bounces with the inclusion of other effects. With the addition of gravity, after the initial rebound and detachment, the droplet undergoes ballistic motion and falls back onto the bath. As noted by Couder et al [13], a way to investigate repeated bouncing behaviour of droplets is to vibrate the liquid bath and achieve a sustained periodic behaviour for a vibrational forcing of sufficient amplitude. Both gravity and vibrational forcing are straightforward additions to the model and a sinusoidal forcing in the gravity term is introduced with the functional form $G(t)=g(1-\Gamma\cos(\omega_{f}t))$. For sufficiently large $\Gamma$ such a forcing generates the Faraday instability: growing standing waves on the free surface known as Faraday waves. The threshold at which Faraday waves are excited is denoted $\Gamma_{F}$. In order for sustained bouncing behaviour to occur, the bath is instead oscillated close to, but below this threshold, that is, for some $\Gamma<\Gamma_{F}$. In what follows we present the results when such a forcing is introduced to the lubrication-mediated model.
@@
Following the methodology presented in Milewski et al [15], we incorporate the forced oscillations of the bath as an oscillating frame of reference. For a given oscillation frequency $\omega_{f}$ and forcing amplitude $\Gamma$, the equations derived in Section 2 for the liquid bath are now given as
@@
and the equation governing the droplet motion $\tilde{Z}(t)$ in the oscillating frame are
@@
The motion in a fixed laboratory frame $Z(t)$ is obtained by the transformation $Z=\tilde{Z}+\Gamma\omega_{f}^{-2}\cos(\omega_{f}t)$. Figure 12 demonstrates the periodic bouncing behaviour obtained with the lubrication-mediated model. To capture these results, the numerical domain was shortened to an integer ratio of the Faraday wavelength of the bath, specifically, corresponding to $L=12$cm. The droplet was then released from a small height, with an initial velocity of $0.1$ms-1. The choice of initial velocity is unimportant as long as the droplet is able to rebound, as the system will eventually stabilise to a periodic state with these parameters.
Figure 12: Periodic droplet rebounds obtained for viscous silicone oil with radius $R=0.83$mm, on a bath oscillated at $80$Hz with forcing amplitude $\Gamma=3$. Plotted with the oscillation period of an undisturbed free surface behind to highlight deviation away from this rest state. The dynamics settles on a periodic pattern corresponding to 4 oscillations of the bath every 2 rebounds of the droplet (denoted as a (4,2) mode), within a period $0.05$ s.
As shown in Figure 12 the stable state is a $(4,2)$ mode [15], which is denoted as such due to the periodic motion requiring 4 oscillations of the bath for every 2 rebounds of the droplet. In the figure, the dashed line indicates the unperturbed surface, as though the behaviour of the droplet were omitted from the system. The upper line then represents the south pole of the drop, and the lower blue line corresponds to the free surface of the liquid bath.
@@
Figure 13: (a) Computed Force-depth plot for L-M model in a (4,2) mode (hence 2 cycles in the figure per period of motion) and fitted data for a linear spring model. (b) Phase-plane dynamics in an inertial frame of reference for the drop and the free-surface. 2 cycles of the (4,2) mode are shown. The drop alternates from a parabolic trajectory (shown together with a dashed line prediction) and an impact state. The free surface alternates between an elliptic trajectory (prediction shown in dashed ellipse in the centre of the figure) and being almost equal to the drop during impact.
In several studies of bouncing drops, a linear spring model has been used to elucidate the effect of impacts, for example off soap films [45] or to describe drops rebounding on a solid hydrophobic surface [46]. A linear model for rebounding off vibrating baths was proposed in the work of Moláček et al. [47], and recently Primkulov et al. [48] used such a model to describe non-resonant effects in pilot wave hydrodynamics. The simplest linear spring model takes the form
$m\frac{d^{2}Z}{dt^{2}}+\gamma\frac{d\bar{Z}}{dt}+\beta\bar{Z}=-mg,\qquad\bar{Z}<0.$ (53)
Here $Z$ is measured in an inertial frame, while $\bar{Z}$ and $\frac{d\bar{Z}}{dt}$ are measured relative to the free surface; $\bar{Z}$ is the depth of impact, and $\frac{d\bar{Z}}{dt}$ is the speed of the drop relative to the free surface speed. These quantities are somewhat ambiguous since the impact itself deforms the surface, but may be approximated as relative to the free surface unperturbed by the impact. The forces exerted by the bath on the droplet are therefore given by
$-\left(\gamma\frac{d\bar{Z}}{dt}+\beta\bar{Z}\right),\qquad\bar{Z}<0.$
[…]
The results are presented in Figure 13, where we show the computed force during impact in our LM model for the simulation in Figure 12. The periodic curves in the $\bar{Z}-F$ plane are composed of two cycles, one for each impact. For each cycle there is a jump in the force at impact, and this elevated force persists (and increases) throughout the downward motion of the drop. During upward motion the force decreases to a much lower level. The area inside each cycle is the net work done on the fluid bath by the droplet, generating the wave motion. An approximate linear fit of the force curve may be made using $\gamma\approx 0.35$ and $\beta\approx 20$ as shown in the figure. While there appear to be some nonlinear effects at play in the dynamics shown, it is plausible that a linear model could provide further simplifications and efficiencies while also capturing essential behaviour. Implementing a linear model including wave generation is beyond the scope of this work.
Figure 13 also shows the corresponding trajectory of our computed LM solution in the inertial $Z-W$ phase plane. The droplet exhibits parabolic flight trajectories and a clear jump in its acceleration at impact, while the free surface has a jump in velocity at the same time. The dashed lines in the figure indicate parabolas matching the flight path of the droplet and an elliptical path marking an undisturbed vertically oscillating free surface (both in a laboratory frame). The figure demonstrates that the free surface returns to this undisturbed oscillation rapidly after droplet lift-off and before the next impact.
6 Conclusion